OpenStudy (anonymous):

A cup of coffee is placed in a room maintaining a constant temperature of 81°F at time t = 0. The temperature of the coffee after t minutes in the room is described by the function Based on this model, how long does it take the coffee to cool to a temperature of 92°F?

OpenStudy (anonymous):

wts the function?

OpenStudy (bahrom7893):

This is easy just plug in 92 instead of C(t) and solve for t

OpenStudy (anonymous):

56.89 minutes

OpenStudy (anonymous):

yeah

OpenStudy (bahrom7893):

92 = 81+101e^(-0.039t) 92-81 = 101e^(-0.039t) 11 = 101e^(-0.039t) 11/101 = e^(-0.039t) 0.1089 = e^(-0.039t) take ln on both sides..

OpenStudy (bahrom7893):

Ln0.1089 = lne^(-0.039t) Ln0.1089 = (-0.039t)*Lne ( Ln(a^b) = b*Ln(a), memorize this ) Ln0.1089 = (-0.039t)*1 ( Ln(e)=1, memorize this ) Ln0.1089 = (-0.039t) t = Ln0.1089 / -0.039

OpenStudy (bahrom7893):

t = 56.85449 according to wolframalpha and yea so roughly 57 hours and hero is right

OpenStudy (anonymous):

Isn't that in minutes?

OpenStudy (bahrom7893):

oh yea i think so.. about an hour haha

OpenStudy (bahrom7893):

*roughly 57 minutes not hours..